# Stationary multiple euclidon solutions to the vacuum Einstein equations

### v1

@@ Abstract
In [1] the non-linear superposition of the stationary euclidon solution with an arbitrary axially symmetric stationary gravitational field on the basis of the method of variation of parameters was constructed. In the paper [1] stationary soliton solution of the Einstein equations was generalized to the case of a stationary seed metric. The formulae obtained in [1], [2] have a simple and compact form, permitting an effective non-linear ”addition” of the solutions. These euclidon solutions serve as ”building blocks” of the theory, which allows for the construction of almost all known solutions to the vacuum static axially-symmetric Einstein equations, including such important ones as the Kerr-NUT solution. The stationary euclidon solution has a clear physical interpretation as a relativistic accelerated non-inertial reference frame, which provides a different perspective on the physical interpretation of well-known solutions, such as the Kerr solution [3].
@@ 1 Introduction
The first exact solutions of the axially symmetric stationary vacuum equations were presented by Lewis [4] and Van Stockum [5]. Later Papapetrou [6] obtained a canonical form of the line element, and Ernst [7] formulated the stationary vacuum problem in the form of a single equation for a complex function. An outstanding result was obtained in 1963 by Kerr [8] , whose solution possibly describes the exterior gravitational field of a rotating source. A family of a asymptotically flat solutions of the stationary vacuum equations was found by Tomimatsu and Sato [9], [10]. Later, a survey of known solutions was given by Kinnersley [11] . Possible generalizations of the Tomimatsu-Sato metric were found by Yamazaki [12], [13] . Hori [14] and Cosgrove [15], [16], and asymptotically not-flat solutions of the Ernst problem were obtained in [17], [19], [4]. The Ernst equation [7] makes it possible to extract the solutions by the method of separation of variables [20], [21] . In a series of papers [22], [23], [24], a non-canonical form of the stationary metrics was used to obtain new solutions of the Einstein equations. An approximate method for constructing solutions depending on two harmonic functions was proposed in [25], [26]. Ernst [27] dreamt of construction of such a kind of solutions, but exact ones.
@@ 2 Basic equations
The operators $\overrightarrow{\nabla}$ and $\Delta$ defined by the formulae
@@
It should be noted that the integrability condition of Equations 2.4, 2.5 for determining the function $\gamma(\rho,z)$ is the system 2.2, 2.3 which does not contain $\gamma(\rho,z)$ Hence, the problem of obtaining exact axially symmetric solutions of the Einstein Equations outside the sources reduces to integration of the differential Equations 2.2, 2.3
In the prolate ellipsoidal coordinates $(x,y)$ the operators $\Delta$ and $\overrightarrow{\nabla}$ take the form
@@
and Equations 2.4, 2.5 take the form
@@
$-2y(1-y^{2})(\frac{\partial f}{\partial x}\frac{\partial f}{\partial y}+\frac{\partial\Phi}{\partial x}\frac{\partial\Phi}{\partial y})\},$ (2.6)
@@
$-2x(1-y^{2})(\frac{\partial f}{\partial x}\frac{\partial f}{\partial y}+\frac{\partial\Phi}{\partial x}\frac{\partial\Phi}{\partial y})\}.$ (2.7)
The line element 2.1 in this case can be rewritten as,
@@
$+(x^{2}-1)(1-y^{2})d\varphi^{2}]-f(dt-\omega d\varphi)^{2},$ (2.8)
@@
There exist two possibilities of writing down Equations 2.2, 2.3 and in a coordinate-free form.
@@ 1. If we switch from the rotation potential $\omega$ to the new potential $\Phi$ by the formulae
This form of the field equations was developed in [28], [29]. In the Weyl canonical coordinates we can rewrite 2.16 as
@@
In this case, using 2.17 and 2.18, we obtain
@@
$\frac{A_{1}}{\partial z}-\frac{A_{2}}{\partial\rho}=0,\quad\frac{B_{1}}{\partial z}-\frac{B_{2}}{\partial\rho}=A_{2}B_{1}-A_{1}B_{2}.$
Let us consider an analogue of the system 2.19 in the ellipsoidal coordinates $(x,y)$.
@@
gives us, from 2.15, the equation
@@
Its proof can be obtained immediately by substitution of into the corresponding Equation in 2.23.
From Equations 2.26 , with the help of 2.22 , one can obtain:
@@
is recognized to be a special case of 2.26 corresponding to the choice $B_{0}=C_{0}=E_{0}=0,\ \ A_{0}=D_{0}$. So 2.27 take the form
@@
As an another particular case of 2.26, the constant phase transformation
@@
where $\alpha_{0}$ is real constant and the function $\tilde{\xi}$ is introduced according 2.24, gives us a possibility of obtaining new solutions of the Ernst equation 2.25 from old ones.
@@
where $C_{1},\ \ C_{2},\ \ C_{3}$ and $C_{4}$ are real constants, preserves the characteristic form of the Papapetrou line element 2.1. In this connection we can proof the following theorem.
@@ 2. The substitution $f=\rho/F$ in equations gives us the following equations:
From 2.37 with the help of 2.35 we obtain
@@
In 1932 Lewis [4] obtained the following class of exact solutions of Equations (ref2.13, 2.15) and:
@@
Here $C_{1}$, $C_{2}$, $C_{3}$ and $U_{0}$ are arbitrary constants, $\psi(\rho,z)=(1/\rho)(\partial\chi/\partial\rho)$ and $\chi(\rho,z)$ are functions satisfying the linear Equations 2.9, 2.11.
Note that Weyl’s static solution $f_{\mathcal{W}}=(\rho/C_{1})e^{-\psi},\Phi_{\mathcal{W}}=C_{2}$ and $\omega_{\mathcal{W}}=C_{3}$ (where $C_{2}$ and $C_{3}$ are usually set equal to zero) can be obtained from 3.1 in the limit $U_{0}\rightarrow\infty.$
To obtain the basic classes of stationary vacuum solutions in a more or less unified way, we shall use the representation of Equations 2.15 in the form of four first-order differential equations 2.19 for the functions 2.18 , which we shall further subject to certain algebraic dependences.
The class of solutions 3.1 can be obtained by making the functions 2.18satisfy the condition
@@
Then from 2.19 we have the equation
@@
where the functions $\psi$ and $\chi$ satisfy Equations 2.9 and 2.11, respectively, and are related to each other by $\psi=(1/\rho)(\partial\chi/\partial\rho).$ The function $\gamma^{\prime}$ can be obtained from Equations.
@@
It should be noted that the solution 3.7 can be transformed to the static vacuum Weyl class by a complex linear transformation of the coordinates $\varphi$ and $t$ ([55]).
@@
The function $\gamma^{\prime}$ can be obtained from 3.8.
@@
It can be checked that a direct substitution of the solution 3.12 nullifies all components of the Riemann tensor, which corresponds to flat space.
@@ 2. Let us finally consider simple methods of solving Equations
The integration 3.19, as well as the following choice and renaming of the constants,
@@
In a special case, setting in 2.38
@@ 4 The euclidon method
$+\left[(z-z_{1})\sinh U+\sqrt{\rho^{2}+(z-z_{1})^{2}}\cdot\cosh U\right]\frac{1}{f^{0}}\frac{\partial\Phi^{0}}{\partial\rho}+$
$+\rho\frac{\sinh U}{f^{0}}\frac{\partial\Phi^{0}}{\partial z},$ (4.5)
@@
$+\left[(z-z_{1})\sinh U+\sqrt{\rho^{2}+(z-z_{1})^{2}}\cdot\cosh U\right]\frac{1}{f^{0}}\frac{\partial\Phi^{0}}{\partial z}-\rho\frac{\sinh U}{f^{0}}\frac{\partial\Phi^{0}}{\partial\rho}.$
@@
$+[\cosh U+(\frac{xy-1}{x-y})\sinh U]\frac{1}{f^{0}}\frac{\partial\Phi^{0}}{\partial x}+(\frac{1-y^{2}}{x-y})\frac{\sinh U}{f^{0}}\frac{\partial\Phi^{0}}{\partial y},$
@@
$+[\cosh U+(\frac{xy-1}{x-y})\sinh U]\frac{1}{f^{0}}\frac{\partial\Phi^{0}}{\partial y}-(\frac{x^{2}-1}{x-y})\frac{\sinh U}{f^{0}}\frac{\partial\Phi^{0}}{\partial x}.$
@@ 1-I. If we choose
in this case we obtain from
@@ 1-II. If we choose
in this case we obtain from
@@
$-\ln[c_{0}(1+y)\sqrt{1+\tanh U_{0}}-(x-1)\sqrt{1-\tanh U_{0}}].$ (5.12)
@@ 6 Superposition of two-stationary euclidon solution with an arbitrary stationary vacuum field
$+[\cosh U^{a}+(\frac{xy+1}{x+y})\sinh U^{a}]\frac{1}{f_{0}}\frac{\partial\Phi_{0}}{\partial x}+(\frac{1-y^{2}}{x+y})\frac{\sinh U^{a}}{f_{0}}\frac{\partial\Phi_{0}}{\partial y},$
@@
$+[\cosh U^{a}+(\frac{xy+1}{x+y})\sinh U^{a}]\frac{1}{f_{0}}\frac{\partial\Phi_{0}}{\partial y}-(\frac{x^{2}-1}{x+y})\frac{\sinh U^{a}}{f_{0}}\frac{\partial\Phi_{0}}{\partial x}.$
@@
$+[\cosh U^{b}+(\frac{xy-1}{x-y})\sinh U^{b}]\frac{1}{f_{0}}\frac{\partial\Phi_{0}}{\partial x}+(\frac{1-y^{2}}{x-y})\frac{\sinh U^{b}}{f_{0}}\frac{\partial\Phi_{0}}{\partial y},$
@@
$+[\cosh U^{b}+(\frac{xy-1}{x-y})\sinh U^{b}]\frac{1}{f_{0}}\frac{\partial\Phi_{0}}{\partial y}-(\frac{x^{2}-1}{x-y})\frac{\sinh U^{b}}{f_{0}}\frac{\partial\Phi_{0}}{\partial x}.$

### v3

@@ Abstract
This paper is dedicated to Professor Ts. I. Gutsunaev, who passed away. In [1], we constructed the non-linear superposition of a stationary euclidon solution with an arbitrary axially symmetric stationary gravitational field using the method of variation of parameters. In the same work [1], the stationary soliton solution of the Einstein equations was generalized to the case of a stationary seed metric. The formulae obtained in [1, 2] have a simple, compact structure that enables an effective non-linear “addition” of solutions. In this framework, euclidon solutions serve as building blocks that generate almost all known vacuum, static, axially symmetric solutions of the Einstein equations, including the important Kerr–NUT family. Moreover, the stationary euclidon solution admits a clear physical interpretation as a relativistically accelerated, non-inertial reference frame, providing an alternative perspective on the physical meaning of celebrated solutions such as Kerr [3].
@@ 1 Introduction
In [1], we constructed the non-linear superposition of a stationary euclidon solution with an arbitrary axially symmetric stationary gravitational field using the method of variation of parameters. In the same work [1], the stationary soliton solution of the Einstein equations was generalized to the case of a stationary seed metric. The formulae obtained in [1, 2] have a simple, compact structure that enables an effective non-linear “addition” of solutions. In this framework, euclidon solutions serve as building blocks that generate almost all known vacuum, static, axially symmetric solutions of the Einstein equations, including the important Kerr–NUT family. Moreover, the stationary euclidon solution admits a clear physical interpretation as a relativistically accelerated, non-inertial reference frame, providing an alternative perspective on the physical meaning of celebrated solutions such as Kerr [3]. The first exact solutions of the axially symmetric stationary vacuum equations were presented by Lewis [4] and Van Stockum [5]. Later Papapetrou [6] obtained a canonical form of the line element, and Ernst [7] formulated the stationary vacuum problem in the form of a single equation for a complex function. An outstanding result was obtained in 1963 by Kerr [8] , whose solution possibly describes the exterior gravitational field of a rotating source. A family of a asymptotically flat solutions of the stationary vacuum equations was found by Tomimatsu and Sato [9], [10]. Later, a survey of known solutions was given by Kinnersley [11] . Possible generalizations of the Tomimatsu-Sato metric were found by Yamazaki [12], [13] . Hori [14] and Cosgrove [15], [16], and asymptotically not-flat solutions of the Ernst problem were obtained in [17], [19], [4]. The Ernst equation [7] makes it possible to extract the solutions by the method of separation of variables [20], [21] . In a series of papers [22], [23], [24], a non-canonical form of the stationary metrics was used to obtain new solutions of the Einstein equations. An approximate method for constructing solutions depending on two harmonic functions was proposed in [25], [26]. Ernst [27] dreamt of construction of such a kind of solutions, but exact ones.
2 Author Declarations
The authors have no conflicts to disclose.
@@ 3 Basic equations
[…]
@@
It should be noted that the integrability condition of Equations 3.4, 3.5 for determining the function $\gamma(\rho,z)$ is the system 3.2, 3.3 which does not contain $\gamma(\rho,z)$ . Hence, the problem of obtaining exact axially symmetric solutions of the Einstein Equations outside the sources reduces to integration of the differential Equations 3.2, 3.3
The prolate ellipsoidal coordinates $(x,y)$ are related to the Weyl canonical coordinates $(\rho,z)$ by the formulae
$x=\frac{1}{2k_{0}}[\sqrt{\rho^{2}+(z+k_{0})^{2}}+\sqrt{\rho^{2}+(z-k_{0})^{2}}],$
$y=\frac{1}{2k_{0}}[\sqrt{\rho^{2}+(z+k_{0})^{2}}-\sqrt{\rho^{2}+(z-k_{0})^{2}}],$
where $k_{0}$ is a real constant.
The inverse transformation is
$\rho=k_{0}\sqrt{(x^{2}-1)(1-y^{2})},\quad z=k_{0}xy.$
[…]
@@
and Equations 3.4, 3.5 take the form
@@
$2y(1-y^{2})(\frac{\partial f}{\partial x}\frac{\partial f}{\partial y}+\frac{\partial\Phi}{\partial x}\frac{\partial\Phi}{\partial y})\},$ (3.6)
@@
$2x(1-y^{2})(\frac{\partial f}{\partial x}\frac{\partial f}{\partial y}+\frac{\partial\Phi}{\partial x}\frac{\partial\Phi}{\partial y})\}.$ (3.7)
The line element 3.1 in this case can be rewritten as,
@@
$(x^{2}-1)(1-y^{2})d\varphi^{2}]-f(dt-\omega d\varphi)^{2},$ (3.8)
@@
There exist two possibilities of writing down Equations 3.2, 3.3 and in a coordinate-free form.
@@ 1. If we switch from the rotation potential $\omega$ to the new potential $\Phi$ by the formulae
This form of the field equations was developed in [28], [29]. In the Weyl canonical coordinates we can rewrite 3.16 as
@@
In this case, using 3.17 and 3.18, we obtain
@@
$\frac{\partial A_{1}}{\partial z}-\frac{\partial A_{2}}{\partial\rho}=0,\quad\frac{\partial B_{1}}{\partial z}-\frac{\partial B_{2}}{\partial\rho}=A_{2}B_{1}-A_{1}B_{2}.$
Let us consider an analogue of the system 3.19 in the ellipsoidal coordinates $(x,y)$.
@@
gives us, from 3.15, the equation
@@
Its proof can be obtained immediately by substitution of into the corresponding Equation in 3.23.
From Equations 3.26 , with the help of 3.22 , one can obtain:
@@
is recognized to be a special case of 3.26 corresponding to the choice $B_{0}=C_{0}=E_{0}=0,\ \ A_{0}=D_{0}$. So 3.27 take the form
@@
As an another particular case of 3.26, the constant phase transformation
@@
where $\alpha_{0}$ is real constant and the function $\tilde{\xi}$ is introduced according 3.24, gives us a possibility of obtaining new solutions of the Ernst equation 3.25 from old ones.
@@
where $C_{1},\ \ C_{2},\ \ C_{3}$ and $C_{4}$ are real constants, preserves the characteristic form of the Papapetrou line element 3.1. In this connection we can proof the following theorem.
@@ 2. The substitution $f=\rho/F$ in equations gives us the following equations:
From 3.37 with the help of 3.35 we obtain
@@
In 1932 Lewis [4] obtained the following class of exact solutions of Equations (ref2.13, 3.15) and:
@@
Here $C_{1}$, $C_{2}$, $C_{3}$ and $U_{0}$ are arbitrary constants, $\psi(\rho,z)=(1/\rho)(\partial\chi/\partial\rho)$ and $\chi(\rho,z)$ are functions satisfying the linear Equations 3.9, 3.11.
Note that Weyl’s static solution $f_{\mathcal{W}}=(\rho/C_{1})e^{-\psi},\Phi_{\mathcal{W}}=C_{2}$ and $\omega_{\mathcal{W}}=C_{3}$ (where $C_{2}$ and $C_{3}$ are usually set equal to zero) can be obtained from 4.1 in the limit $U_{0}\rightarrow\infty.$
To obtain the basic classes of stationary vacuum solutions in a more or less unified way, we shall use the representation of Equations 3.15 in the form of four first-order differential equations 3.19 for the functions 3.18 , which we shall further subject to certain algebraic dependences.
The class of solutions 4.1 can be obtained by making the functions 3.18satisfy the condition
@@
Then from 3.19 we have the equation
@@
where the functions $\psi$ and $\chi$ satisfy Equations 3.9 and 3.11, respectively, and are related to each other by $\psi=(1/\rho)(\partial\chi/\partial\rho).$ The function $\gamma^{\prime}$ can be obtained from Equations.
@@
It should be noted that the solution 4.7 can be transformed to the static vacuum Weyl class by a complex linear transformation of the coordinates $\varphi$ and $t$ ([55]).
@@
The function $\gamma^{\prime}$ can be obtained from 4.8.
@@
It can be checked that a direct substitution of the solution 4.12 nullifies all components of the Riemann tensor, which corresponds to flat space.
@@ 2. Let us finally consider simple methods of solving Equations
The integration 4.19, as well as the following choice and renaming of the constants,
@@
In a special case, setting in 3.38
@@ 5 The euclidon method
$\left[(z-z_{1})\sinh U+\sqrt{\rho^{2}+(z-z_{1})^{2}}\cdot\cosh U\right]\frac{1}{f^{0}}\frac{\partial\Phi^{0}}{\partial\rho}+$
$\rho\frac{\sinh U}{f^{0}}\frac{\partial\Phi^{0}}{\partial z},$ (5.5)
@@
$\left[(z-z_{1})\sinh U+\sqrt{\rho^{2}+(z-z_{1})^{2}}\cdot\cosh U\right]\frac{1}{f^{0}}\frac{\partial\Phi^{0}}{\partial z}-\rho\frac{\sinh U}{f^{0}}\frac{\partial\Phi^{0}}{\partial\rho}.$
@@
$\cosh U+(\frac{xy-1}{x-y})\sinh U]\frac{1}{f^{0}}\frac{\partial\Phi^{0}}{\partial x}+(\frac{1-y^{2}}{x-y})\frac{\sinh U}{f^{0}}\frac{\partial\Phi^{0}}{\partial y},$
@@
$\cosh U+(\frac{xy-1}{x-y})\sinh U]\frac{1}{f^{0}}\frac{\partial\Phi^{0}}{\partial y}-(\frac{x^{2}-1}{x-y})\frac{\sinh U}{f^{0}}\frac{\partial\Phi^{0}}{\partial x}.$
@@ 1-I. If we choose
in this case we obtain from 6.3, 6.4
@@ 1-II. If we choose
in this case we obtain from 6.3, 6.4
@@
$\ln[c_{0}(1+y)\sqrt{1+\tanh U_{0}}-(x-1)\sqrt{1-\tanh U_{0}}].$ (6.12)
@@
If we put
$\frac{1+a_{0}b_{0}}{\sqrt{(1+a^{2}_{0})(1+b^{2}_{0})}}=k_{0}(m^{2}_{0}+l^{2}_{0})^{-1/2},\frac{b_{0}-a_{0}}{\sqrt{(1+a^{2}_{0})(1+b^{2}_{0})}}=a(m^{2}_{0}+l^{2}_{0})^{-1/2},$
$\frac{1-a_{0}b_{0}}{\sqrt{(1+a^{2}_{0})(1+b^{2}_{0})}}=m_{0}(m^{2}_{0}+l^{2}_{0})^{-1/2},\frac{b_{0}+a_{0}}{\sqrt{(1+a^{2}_{0})(1+b^{2}_{0})}}=-l_{0}(m^{2}_{0}+l^{2}_{0})^{-1/2},$
we can obtain the Kerr-NUT solution in Boyer-Lindquist coordinates $(r,\theta)\ \ (x=(r-m_{0})/k_{0},\ \ y=cos\theta)$
$ds^{2}=\frac{r^{2}+(a\cos\theta-l_{0})^{2}}{r^{2}-2m_{0}r+a^{2}-l^{2}_{0}}dr^{2}+[r^{2}+(a\cos\theta-l_{0})^{2}]\times$
$\times[d\theta^{2}+\frac{(r^{2}-2m_{0}r+a^{2}-l^{2}_{0})\sin^{2}\theta d\varphi^{2}}{r^{2}-2m_{0}r+a^{2}\cos^{2}\theta-l^{2}_{0}}]-[1-2\frac{m_{0}r+l_{0}(l_{0}-a\cos\theta)}{r^{2}+(a\cos\theta-l_{0})^{2}}]\times$
$\times\{dt-(\frac{2a\sin^{2}\theta[m_{0}r+l_{0}(l_{0}-a\cos\theta)]}{r^{2}-2m_{0}r+a^{2}\cos^{2}\theta-l^{2}_{0}}-2l_{0}\cos\theta)d\varphi\}^{2}.$
@@ 7 Superposition of two-stationary euclidon solution with an arbitrary stationary vacuum field
$\cosh U^{a}+(\frac{xy+1}{x+y})\sinh U^{a}]\frac{1}{f_{0}}\frac{\partial\Phi_{0}}{\partial x}+(\frac{1-y^{2}}{x+y})\frac{\sinh U^{a}}{f_{0}}\frac{\partial\Phi_{0}}{\partial y},$
@@
$\cosh U^{a}+(\frac{xy+1}{x+y})\sinh U^{a}]\frac{1}{f_{0}}\frac{\partial\Phi_{0}}{\partial y}-(\frac{x^{2}-1}{x+y})\frac{\sinh U^{a}}{f_{0}}\frac{\partial\Phi_{0}}{\partial x}.$
@@
$\cosh U^{b}+(\frac{xy-1}{x-y})\sinh U^{b}]\frac{1}{f_{0}}\frac{\partial\Phi_{0}}{\partial x}+(\frac{1-y^{2}}{x-y})\frac{\sinh U^{b}}{f_{0}}\frac{\partial\Phi_{0}}{\partial y},$
@@
$\cosh U^{b}+(\frac{xy-1}{x-y})\sinh U^{b}]\frac{1}{f_{0}}\frac{\partial\Phi_{0}}{\partial y}-(\frac{x^{2}-1}{x-y})\frac{\sinh U^{b}}{f_{0}}\frac{\partial\Phi_{0}}{\partial x}.$
In this way, many generalizations of the Kerr solution describing axially symmetric rotating masses have been obtained [1].